OpenStudy (anonymous):

if something is invested at 6% quarterly, would i multiply 4 to 0.06 of divide 4 by 0.06?

OpenStudy (anonymous):

That's compound interest - after the 1st quarter you earn 6% of 100, but after the second quarter you'll earn 6% of 106, and so on.

OpenStudy (anonymous):

ye.. its asking me how long it will take for $500 to triple if it is invested at 6% quarterly..

OpenStudy (anonymous):

You can use the formula for compound interest, or work it out by hand if it's a small amount. If it's academic I think we'll want to do it the fancy way

OpenStudy (anonymous):

So we use final amount = initial amount * (1 + interest rate) ^ n

OpenStudy (anonymous):

where n is how many times you add the interest

OpenStudy (anonymous):

yee..

OpenStudy (anonymous):

in this case we're looking for n

OpenStudy (anonymous):

Yep. Do you know how to get n from there?

OpenStudy (anonymous):

well okay so, this is what i did. A=P(1+r)^t 1500=500(1+0.15)^t 3=(1.015)^t ln3=tln(1.015) ln3/ln1.015=t 73.79=t

OpenStudy (anonymous):

We can't divide the interest rate by 4. Just use 6%, and then divide t (which will be in 3-month periods) by 4 to get the answer in years.

OpenStudy (anonymous):

By the way I rather like the giraffe avatar. I gotta get myself some giraffes of my own.

OpenStudy (anonymous):

@jirafa oooh! okay so you're saying i would have to do A=P(1+r)^t 1500=500(1+0.06)^t/4 3=(1.06)^t/4 ln3=(t/4)ln(1.015) 4ln3/ln1.015=t

OpenStudy (anonymous):

lol and yes yes thank you :)

OpenStudy (anonymous):

oh my bad... i meant ln3=(t/4)ln(1.06) 4ln3/ln1.06=t

OpenStudy (anonymous):

?:S

OpenStudy (anonymous):

There you go!

OpenStudy (anonymous):

or would i divide it by 4 after i get t?

OpenStudy (anonymous):

Doesn't matter

OpenStudy (anonymous):

No, hold on a second

OpenStudy (anonymous):

You can either divide by 4 after getting it, or use 4t in the equation

OpenStudy (anonymous):

Using t/4 in the equation will convert from years to three month periods, the opposite of what we want

OpenStudy (anonymous):

so i would have to work with 1500 = 500 (1.06)^4t

OpenStudy (anonymous):

Yep, that's it. I take it that the interest is 6% quarterly, not 6% yearly but compounded quarterly?

OpenStudy (anonymous):

If the latter, you need another formula

OpenStudy (anonymous):

@Chlorophyll oooh! that works!

OpenStudy (anonymous):

@jirafa & @Chlorophyll thanks for your help!

OpenStudy (anonymous):

@Chlorophyll so i would do the same if it were compounded daily right?

OpenStudy (anonymous):

Yeah, there you go. Should've assumed that the interest rate was annual, that's what everyone and their cat does anyway

OpenStudy (anonymous):

Cheers

OpenStudy (anonymous):

@jirafa their cats? :S lol

OpenStudy (anonymous):

Isn't that a somewhat common English expression? Well nevermind that

OpenStudy (anonymous):

Cats certainly can't do compound interest problems, they aren't as smart as giraffes

OpenStudy (anonymous):

Later

OpenStudy (anonymous):

woah! wait... when i do the daily one... it gives me a huge decimal.. but its like 1.643835616E-4

OpenStudy (anonymous):

try it out again, it should work all right

OpenStudy (anonymous):

jsut notice that you have to input t as years

OpenStudy (anonymous):

not as days

OpenStudy (anonymous):

Figured it out already?

OpenStudy (anonymous):

yes i did! thankkss :)

OpenStudy (anonymous):

how would i solve: at what rate of interes compounded annually should you invest $500 if you want to have $25000 in 12 years?

OpenStudy (anonymous):

@jirafa

OpenStudy (anonymous):

Same equation, but this time the variable you need is r

OpenStudy (anonymous):

okay. so i got until ln50=12 ln(1+r)

OpenStudy (anonymous):

A=P(1+r)^t 25000=500(1+r)^12 50=(1+r)^12 ln50 = 12 ln (1+r)

OpenStudy (anonymous):

Doing alright so far. ln 50 is a number, so calculate that one and divide both sides.

OpenStudy (anonymous):

its 3.912=12ln(1+r) so do i divide3.912 by 12?

OpenStudy (anonymous):

I'm unsure if you're correct up to this step: 50=(1+r)^12 If you are, there's no need to take ln!

OpenStudy (anonymous):

Yep, there's no unknown in the exponent this time, so we can get rid of it without using ln at all. How would you remove that 12?

OpenStudy (anonymous):

um... do ((1+r)^12)^1/12 ? :S

OpenStudy (anonymous):

That'll get rid of the 12 all right, but you have to do it in both sides. x^(1/12) = \[\sqrt[12]{x}\] isn't it?

OpenStudy (anonymous):

oh okay so \[\sqrt[12]{50}=\sqrt[12]{(1+r)^{12}}\]

OpenStudy (anonymous):

? :s

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

YAY! okay thanks so much for your help again @jirafa @Chlorophyll i shall go to bed now.. its getting late

OpenStudy (anonymous):

i got 38.5% btw.

OpenStudy (anonymous):

The rate is 38.5 % :0 That's ridiculous rip off !!!

OpenStudy (anonymous):

Hah, you definitely need some hardcore interest rate to make 25 grand from $500. By all means do tell me which bank offered you this.

OpenStudy (anonymous):

Night